Motorcycle travels 45/mph, walk 6/mph, if travel 25 miles in 2 hrs how far did motorcycle go before broke down?
Have m45/60+w6/60=25/2
=m45+w6=750=w6=750-m45=w=12.5-6.5m so
but onot sure of that and don't know how to formulate rest of problem.

Found 2 solutions by mananth, mtdddog75:
Answer by mananth(16949)

Motorcycle travels 45/mph, walk 6/mph, if travel 25 miles in 2 hrs how far did motorcycle go before broke down?

let motorcycle be driven for x kms.
walking will be 25-x
distance/rate = time
x/45 is time driven
25-x/6 = time walked
x/45 + 25-x / 6= 2 270 LCD
6x+1125-45x = 270*2
-39x=540-1125
-39x=-585
x= -585 / -39
x=15 the motor cycle broke down after 15 kilometers
& he walked 10 kilometers







Answer by mtdddog75(2)


x should be time traveled hours
45x motorcycle motorcycle distance
6(x-2) + 25 walking distance

45x = 6(x-2) + 25 starting eq.
6x - 12 + 25 get rid of par. and combine like terms
45x = 6x + 13
-6x move variable to left side
39x = 13
39x/39 <= 13/39
divide by 39 to leave x on left
x=1/3hr = 20 minutes motorcycle broke down after 20 minutes 1/3 of 45 is 15miles + 1 hour 40 minutes of walking 10 miles = 25 miles
Mike